# MorphSeek: Fine-grained Latent Representation-Level Policy Optimization for Deformable Image Registration

### v1

@@ 4.1 Datasets
We evaluate MorphSeek on three registration tasks spanning five 3D medical imaging datasets.
OASIS. A Learn2Reg[1601learn2reg] task for inter-patient brain MRI registration[1401]. Volumes are preprocessed and resampled to $160\times 192\times 224$. From 414 training volumes, we form 400/100/20 pairs for pretraining/GRPO/validation. Test set: 19 official challenge validation pairs.
LiTS. Contains 131 contrast-enhanced abdominal CT scans[1403]. We use whole-liver labels only. Volumes are resampled to $160\times 192\times 224$. Data split: 400/100/20/40 pairs for pretraining/GRPO/validation/test, with no test volume overlap in training.
Abdomen MR-CT. A Learn2Reg[1601learn2reg] task for intra-patient abdominal MR-CT registration using 8 paired scans from TCIA[14021] and 90 unpaired scans from BCV[14022] and CHAOS[14023]. Volumes are resampled to $160\times 192\times 192$. From unpaired data, we construct 400/100/20 MR-CT pairs for pretraining/GRPO/validation. Test set: 8 official paired scans. For this cross-modality task we use the MIND descriptor[0808mindmetric] instead of MSE as the image similarity term.
@@ 4.2 Baselines and Implementations
We select three baseline algorithms for refactoring, namely VoxelMorph[0501voxelmorph], TransMorph[0502transmorph], and NICE-Trans[0504nicetrans], and we include two state-of-the-art frameworks—CorrMLP[1501corrMLP], RIIR[0904RIIRstep]—and an RL-based framework SPAC[0801SPACweibiaoqian], for comparison. To ensure fairness, each baseline is trained with the same weakly supervised setting and a Dice loss on segmentation labels. We also include WarpDDF+RegCut[1201semi] which can combine segmentation labeled data and unlabeled data, as a comparison. Because most traditional optimization-based methods cannot effectively leverage label supervision under the same setting, we place their unsupervised training results in the supplementary materials.
We additionally adopt a widened VoxelMorph variant (VoxelMorph-L) with channels [32, 64, 128, 256, 256] (approximately 27M parameters) to provide a sufficiently expressive latent space for GRPO. We also use VoxelMorph-L as the baseline.
Training uses Adam (1e-4) and batch size 1; other hyper-parameters and hardware details are in the supplement.
We report Dice[1701dice] (%) for segmentation overlap and the percentage of voxels with negative Jacobian determinant (NJD[1703njd], %) for deformation regularity.
@@ 8 OOM Error N/A N/A N/A
# Configuration
Sample
encoder $\mathbf{f_{L}}$?
Weak
Supervision?
Step/
Traj
VoxelMorph-L TransMorph NICE-Trans
@@ 5.3 Computational Overhead and the Role of Warm-up
Table 4 analyzes the structural and runtime overhead introduced by the RL-friendly refactoring. For all three backbones, the additional parameters are below 3% and single-step inference latency remains almost unchanged. Multi-step inference exhibits roughly linear growth in runtime (about threefold latency for three refinement steps), which is expected for a cascaded scheme and enables a simple trade-off between accuracy and computational budget at deployment time.
Table 4: Efficiency analysis of RL-friendly refactoring on OASIS task. Inference time: averaged over five runs on the test set.
Baseline Model Parameters GPU Inference Time (ms)
Original $\Delta$ Abs $\Delta$ Rel Original +MorphSeek 1/2/3 step(s)
VoxelMorph-L 27.05M +0.13M +0.48% 625 685 / 1387 / 2022
TransMorph 46.77M +1.18M +2.53% 401 444 / 900 / 1376
NICE-Trans 5.71M +0.13M +2.27% 406 431 / 864 / 1295
The training curves on OASIS (Figure 5) highlight the critical role of unsupervised warm-up. Without warm-up, GRPO training is prone to oscillating policy gradients, higher sensitivity to hyperparameters, and a greater risk of non-physical deformations. With warm-up, the model typically reaches the same target Dice in roughly half the fine-tuning epochs and exhibits much smoother validation curves.
Figure 5: Validation Dice on OASIS: Effect of Warm-up Before GRPO (TransMorph Backbone), where the dotted line marks the switch from warm-up to GRPO
We therefore position warm-up as a prior-shaping and cost-reduction stage: it does not necessarily raise the ultimate performance ceiling, but substantially reduces the time, computational resources, and instability risks required to reach a given accuracy level, by pre-aligning the latent space before policy optimization.

### v3

@@ 4.1 Datasets
[…]
@@ 4.2 Baselines and Implementations
We refactor VoxelMorph-L, TransMorph, and NICE-Trans under the same weakly supervised setting, using identical pair lists, the same labeled pairs, and comparable epoch budgets. We further compare against CorrMLP[36], RIIR[53], SPAC[28], and WarpDDF+RegCut[25]; for methods whose public code or training recipes are not directly compatible with our unified setting, we follow their published protocols. VoxelMorph-L uses channels [32, 64, 128, 256, 256] to provide a sufficiently expressive latent space for GRPO. Training uses Adam (1e-4) with batch size 1; detailed hyperparameters and hardware are given in the supplement.
We report Dice[10] (%) for segmentation overlap and the percentage of voxels with negative Jacobian determinant (NJD[24], %) for deformation regularity.
@@ 5 Results and Analysis
Figure 2: Representative visual comparisons across the three registration tasks. Labels are overlaid only for the two abdominal tasks; OASIS is left unlabeled to avoid clutter from its 35 foreground classes. Additional visual results are provided in the supplement.
@@ 8 OOM Error N/A N/A N/A
# Configuration Sample encoder $\mathbf{f_{L}}$? Weak Supervision? Step/ Traj VoxelMorph-L TransMorph NICE-Trans
@@ 5.3 Computational Overhead and the Role of Warm-up
MorphSeek introduces less than 3% additional parameters across all three backbones, and single-step inference latency remains close to the original models. Multi-step inference scales approximately linearly with the number of refinement steps, providing a simple deployment-time trade-off between accuracy and runtime; detailed measurements are reported in the supplement.
The training curves on OASIS (Figure 4) highlight the critical role of unsupervised warm-up. The blue curve denotes unsupervised warm-up, whereas the green curve denotes a supervised Dice baseline using 100 labeled pairs; the dotted line marks the switch from warm-up to GRPO, so the two curves are not expected to coincide before policy optimization starts. Without warm-up, GRPO training is prone to oscillating policy gradients, higher sensitivity to hyperparameters, and a greater risk of non-physical deformations. In 20 independent runs (TransMorph backbone), warm-up increases stable-training success rate from 33% to 79% and reduces the average convergence epoch from approximately 120 to 75, while also producing smoother validation curves. Supplementary posterior-collapse analysis further shows that unstable runs can become nearly deterministic: sampling the same input ten times yields near-zero Dice variance after collapse, whereas normal warm-started checkpoints retain non-trivial output variance. This confirms that warm-up not only accelerates convergence but also preserves the stochastic exploration capacity required by GRPO.
Figure 4: Validation Dice on OASIS with the TransMorph backbone. The blue curve denotes unsupervised warm-up, the green curve denotes a supervised Dice baseline using 100 labeled pairs, and the dotted line marks the switch from warm-up to GRPO.
We therefore position warm-up as a prior-shaping and cost-reduction stage: it does not necessarily raise the ultimate performance ceiling, but substantially reduces the time, computational resources, and instability risks required to reach a given accuracy level, by pre-aligning the latent space before policy optimization. Supplementary failure cases further show that removing the similarity term from $\mathcal{L}_{\text{warm}}$ can yield seemingly competitive proxy metrics yet visibly smeared anatomy, reinforcing its role as an anatomy-preserving prior.